### 04 Mar 2026 v1

# spacing=nonfrench

Abstract.
To a height-$1$ formal group defined over the ring of integers of a finite extension $K$ of $\mathbb{Q}_{p}$ is attached its $p$-adic Tate module, which is a crystalline character of $\operatorname{Gal}(\overline{K}/K)$ of Hodge-Tate weight $1$. This association is an equivalence. We prove, over extensions whose ramification index is relatively prime to $p^{2}-p$, that the set of consistent sequences attached to a height-$1$ commuting pair $(f,u)$ of noninvertible and invertible formal power series is a crystalline character of weight $1$, for which $f$ is an endomorphism. As a result, we deduce a proof of a conjecture of Lubin in new cases.
1991 Mathematics Subject Classification:
11S82, 11S31, 37P20; 14L05, 11F80, 11F85
The author was supported by the FRIA of the Fonds de la Recherche Scientifique – FNRS
Introduction
Pairs of analytic transformations of a $p$-adic open unit disk with a fixed point at 0 that commute under composition were studied by Lubin in [Lub94] under the name of $p$-adic dynamical systems. He pointed out in ibidem that “for an invertible series to commute with a noninvertible series, there must be a formal group somehow in the background”. This observation has initiated a wave of results in this direction, see for instance [Li96], [Li02], [Sar05], [Sar10], [SS13], [Ber19]. More recently, $p$-adic Hodge theory has proved to be an efficient way for tackling this problem, as in [Ber16], [Spe18], [Poy24], and this paper is written in that vein.
Lubin’s conjecture ([Sar05, p. 131]).
Suppose that $f$ and $u$ are a noninvertible and a nontorsion invertible series, respectively, defined over the ring of integers $\mathcal{O}$ of a finite extension of $\mathbb{Q}_{p}$. Suppose further that the roots of $f$ and all of its iterates are simple, and that $f^{\prime}(0)$ is a uniformizer in $\mathbb{Z}_{p}$. If $f\circ u=u\circ f$, then $f,u\in\operatorname{End}_{\mathcal{O}}(F)$ for some formal group $F$ over $\mathcal{O}$.
Throughout this article, we refer to the above statement as Lubin’s conjecture. The height of the dynamical system together with its ring of definition is a common way to decompose this conjecture into cases. The height-$1$ case over $\mathbb{Z}_{p}$ is solved by Specter in [Spe18] where he recovered the underlying Lubin-Tate power series. Shortly after, Berger solved in [Ber19] all the cases of finite height over $\mathbb{Z}_{p}$ using only $p$-adic analysis. In this article, we solve the height-$1$ cases over extensions whose ramification index is relatively prime to $p^{2}-p$, provided that the linear coefficient of the invertible series lies in $\mathbb{Z}_{p}$. The remaining cases of Lubin’s conjecture are still open nowadays. Our approach consists in generalizing Specter’s analysis of the set of consistent sequences of the dynamical system which yields a Galois character, recovering the missing $\mathbb{Z}_{p}$-module structure on it, and, from this data, reconstructing the latent formal group via integral $p$-adic Hodge-theoretic methods.
Notations
Let $p$ be any prime number and fix an algebraic closure of $\mathbb{Q}_{p}$ containing all our extensions. Define $\mathbb{C}_{p}$ to be its $p$-adic completion, with open unit disk $\mathfrak{m}_{\mathbb{C}_{p}}\!$. If $E$ is a finite extension of $\mathbb{Q}_{p}$, we let $\mathcal{O}_{E}$ be its ring of integers, with maximal ideal $\mathfrak{m}_{E}$, and we let $e_{E}$ be the ramification index of the extension, or $e$ if it is unambiguous. Given a ring $R$, we define $\mathscr{S}_{0}(R)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=XR[\![X]\!]$ to be the set of formal power series over $R$ without constant term. Composition induces a law on $\mathscr{S}_{0}(R)$ making it into a noncommutative monoid, and a necessary and sufficient condition on $s\in\mathscr{S}_{0}(R)$ to be invertible is that $s^{\prime}(0)$ is a unit of $R$. We deal exclusively with rings $R$ arising from a $p$-adic situation, that is, $p$-adic fields, their integers and finite fields of characteristic $p$.
Motivation and main statement
Before stating our main theorem, we remind the reader of some elementary facts about formal groups, see also [Haz12] for an extended presentation. By a formal group, we mean a one-dimensional commutative formal group law. Given a formal group $F$ defined over the ring of integers of a finite extension $K/\mathbb{Q}_{p}$, its ring of $\mathcal{O}_{K}$-endomorphisms, namely $\operatorname{End}_{\mathcal{O}_{K}}(F)$, is a composition-commuting family of noninvertible and invertible formal power series, provided with an embedding of $\mathbb{Z}_{p}$ sending $m\in\mathbb{Z}_{p}$ to $m]_{F}$, the multiplication-by-$m$ map in $F$. The map taking the linear coefficient in $\operatorname{End}_{\mathcal{O}_{K}}(F)$ is injective. If we further assume that $F$ is of height $1$ and $K$ is a nontrivial extension of $\mathbb{Q}_{p}$, then $F$ is not Lubin-Tate and this is the major difference with Specter’s case. What remains true is that $\operatorname{End}_{\mathcal{O}_{K}}(F)$ is a free $\mathbb{Z}_{p}$-algebra of rank $1$, implying that the linear coefficient of every $\mathcal{O}_{K}$-endomorphism of $F$ lies in $\mathbb{Z}_{p}$. We will refer to this fact as (1). An important data attached to the formal group $F$ is its $p$-adic Tate module
[…]
where $\Lambda_{n}(F)$ is the group of $p^{n}$-torsion points of $F$ in $\mathfrak{m}_{\mathbb{C}_{p}}\!$ for every integer $n\geqslant 0$. This is a crystalline character of the absolute Galois group $G_{K}$ of $K$ of Hodge-Tate weight $1$, meaning that $\operatorname{T}_{p}(F)$ is a free $\mathbb{Z}_{p}$-module of rank $1$ provided with a compatible action of $G_{K}$ that is crystalline of weight $1$. According to [SW13, Corollary 6.2.3], what makes it relevant is that the functor
$\operatorname{T}_{p}\colon\;$Formal Groups of Height 11 $\longrightarrow$ Crystalline Characters of $G_{K}$
  $\operatorname{T}_{p}\colon\;$over $\mathcal{O}_{K}$ of Hodge-Tate Weight 11
is an equivalence, given that those of weight $1$ correspond to connected $p$-divisible groups, that is, formal groups [Tat67, Proposition 1]. Sufficiently explicit functors to recover a formal group from an object on the right-hand side can be extracted from the work of Kisin [Kis06], [Kis09] and Zink [Zin01], [Zin02] in integral $p$-adic Hodge theory.
This equivalence plays a key role in this article. Motivated by the above discussion, and given a power series $f\in\mathscr{S}_{0}(\mathcal{O}_{K})$ that has exactly $p$ roots in the open unit disk of $\mathbb{C}_{p}$, whose iterates have only simple roots and such that $f^{\prime}(0)$ is a uniformizer in $\mathbb{Z}_{p}$, we consider the $G_{K}$-set
[…]
of $f$-consistent sequences of elements in $\mathfrak{m}_{\mathbb{C}_{p}}\!$ (see [Lub94, p. 329]), where $\Lambda_{n}(f)$ is now the set of roots of $f^{\circ n}$ in $\mathfrak{m}_{\mathbb{C}_{p}}\!$ for all integers $n\geqslant 0$. If Lubin’s conjecture holds, the latter should be the $p$-adic Tate module of an integral formal group once $f$ commutes with a nontorsion invertible $u\in\mathscr{S}_{0}(\mathcal{O}_{K})$. Therefore, there should exist a compatible $\mathbb{Z}_{p}$-module structure on $T_{f}$ turning it into a crystalline character of weight $1$ and admitting $f$ as an endomorphism. This is what we achieve to recover, for a restriction of $G_{K}$ from $K$ to a finite Galois extension $L$, when $(e_{K},p^{2}-p)=1$. Based on (1), we assume as well that $u^{\prime}(0)\in\mathbb{Z}_{p}^{\times}$. Although this assumption is reasonable, the author regrets not knowing if it can be deduced from the original hypotheses. The $\mathbb{Z}_{p}[G_{L}]$-module structure on $T_{f}$ allows to recover a formal group over $\mathcal{O}_{L}$ for which $(f,u)$ is a pair of endomorphisms, and the uniqueness of the latent formal group immediately implies the descent to $\mathcal{O}_{K}$.
Main theorem (Theorem 3.1).
Let $K/\mathbb{Q}_{p}$ be a finite extension satisfying $(e_{K},p^{2}-p)=1$. Let $(f,u)$ be a commuting pair of elements in $\mathscr{S}_{0}(\mathcal{O}_{K})$ such that $f$ has exactly $p$ roots in $\mathfrak{m}_{\mathbb{C}_{p}}\!$ and all roots of its iterates are simple, with $f^{\prime}(0)$ a uniformizer in $\mathbb{Z}_{p}\!$ and $u^{\prime}(0)\in\mathbb{Z}_{p}^{\times}$ nontorsion. Then there exists a formal group $F$ over $\mathcal{O}_{K}$ such that $f,u\in\operatorname{End}_{\mathcal{O}_{K}}(F)$.
The case of small ramification (e<p1e<p-1)
Our work has a nontrivial intersection with the one of Berger in [Ber19]. If we assume $(f,u)$ to be a commuting pair over $\mathcal{O}_{K}$ as in our main theorem with $(e,p^{2}-p)=1$ and that e<p1e<p-1, then the same arguments as for [Ber19, Corollary A] can be repeated to show that $f$ and $u$ are endomorphisms of a formal group defined over $\mathcal{O}_{K}$. We do not claim any responsibility in this result since it is just an observation.
Structure of the article
In the first section, we recall some invariants and relations attendant to a $p$-adic dynamical system $(f,u)$ from [Lub94]. Using the main theorem of [Li96], we determine the decreasing part of the Newton polygon of any iterate of $f$. This yields information on the roots of the iterates of $f$ and also on the periodic points of $u$. The situation slightly differs from [Spe18] but a similar path can be followed modulo some new adjustments and arguments. The second section is devoted to the recovery of the Tate module structure on the set $T_{f}$ attached to $(f,u)$. In §2.1, we mimic as in [Spe18] the Galois action on $T_{f}$ by means of the action of the $\mathbb{Z}_{p}$-iterates of $u$, in order to attach a character $\chi_{f}\colon G_{L}\rightarrow\mathbb{Z}_{p}^{\times}$ to the pair $(f,u)$, where $L$ is the finite extension obtained by adjoining to $K$ the fixed points of $u$. In §2.2 and §2.3, we expose the material used in §2.4 to construct a crystalline period of $\chi_{f}$. We show that it has weight $1$. This property implies the existence of a formal group $H$ over $\mathcal{O}_{L}$ satisfying $\operatorname{T}_{p}(H)\simeq\mathbb{Z}_{p}(\chi_{f})$. Although $H$ need not be latent to our dynamical system, it should be isomorphic to it. Therefore, we rely on $H$ as a cornerstone of our constructions. In §2.5, we begin by identifying the counterpart of $(f,u)$ in $H$. This allows us, in a second time, to perform a transport of structure from $\operatorname{T}_{p}(H)$ to $T_{f}$ and turn $T_{f}$ into a $\mathbb{Z}_{p}[G_{L}]$-module. In the third section, we move backward until we get a formal group. From [Kis06], we consider the associated Breuil-Kisin module in §3.1, extend it into a window in §3.2 as in [Kis09], and transform it into a connected $p$-divisible group in §3.3 as in [Zin01]. At every stage, we track the behavior of $f$ on the new object, allowing us to conclude in §3.4 that the formal group associated to the connected $p$-divisible group is the latent one.
Acknowledgments
I am sincerely grateful to my advisor, Prof. Maja Volkov, for her guidance and invaluable advice, to Giovanni Bosco for helpful conversations on $p$-adic Hodge theory, and to Prof. Laurent Berger and Prof. Jonathan Lubin for their support and help.
1. The dynamical system
Let $K$ be a finite extension of $\mathbb{Q}_{p}$ satisfying $(e_{K},p^{2}-p)=1$. Throughout this article, we assume $(f,u)$ to be a commuting pair of formal power series in $\mathscr{S}_{0}(\mathcal{O}_{K})$ such that $f$ has exactly $p$ roots in $\mathfrak{m}_{\mathbb{C}_{p}}\!$ and all roots of its iterates are simple, with $f^{\prime}(0)$ a uniformizer in $\mathbb{Z}_{p}$ and $u^{\prime}(0)\in\mathbb{Z}_{p}^{\times}$ nontorsion. Up to replacing $u$ by one of its iterates, we may assume that $u^{\prime}(0)\in 1+p\mathbb{Z}_{p}$. In this section, we review the immediate data at hand in such a dynamical system which will play a crucial role in the subsequent work.
1.1. Logarithm
In his seminal article [Lub94], Lubin showed that in the presence of a formal power series $s\in\mathscr{S}_{0}(\mathcal{O}_{K})$ such that $s^{\prime}(0)$ is nonzero nor a root of unity, there is a unique invertible power series
$\operatorname{Log}_{s}\in\mathscr{S}_{0}(K)$
satisfying $\operatorname{Log}_{s}^{\prime}(0)=1$ and $\operatorname{Log}_{s}(s(X))=s^{\prime}(0)\operatorname{Log}_{s}(X)$, called the logarithm of $s$. Every power series satisfying the same condition as above and commuting with $s$ has the same logarithm. Denote by $\operatorname{Log}_{f}$ the logarithm of $(f,u)$; it converges on $\mathfrak{m}_{\mathbb{C}_{p}}\!$. If $f$ and $u$ are endomorphisms of an integral formal group, then $\operatorname{Log}_{f}$ is the logarithm of this formal group and it is well known that its derivative has coefficients in $\mathcal{O}_{K}$. It turns out that this is true even for $p$-adic dynamical systems [Ber19, Lemma 2.2]. The conjugation
$F(X,Y)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{Log}_{f}^{\circ-1}(\operatorname{Log}_{f}(X)+\operatorname{Log}_{f}(Y))$
is the unique formal group for which $f$, $u$ are endomorphisms, but it is defined a priori over $K$ rather than $\mathcal{O}_{K}$. Therefore, solving Lubin’s conjecture amounts to show that $F$ is defined over $\mathcal{O}_{K}$. For every $m\in\mathbb{Z}_{p}$, we let
$m]_{f}(X)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{Log}_{f}^{\circ-1}(m\operatorname{Log}_{f}(X))\in\mathscr{S}_{0}(K)$
be the unique power series over $K$ commuting with $f$ and acting on the tangent space at 0 by $m$, which is nothing but the multiplication-by-$m$ map in $F$.
1.2. Torsion points
Whenever we talk about the roots or the fixed points of a power series, we mean its roots or its fixed points, respectively, in the $p$-adic open unit disk $\mathfrak{m}_{\mathbb{C}_{p}}\!$ of $\mathbb{C}_{p}$. A noninvertible formal power series $s_{1}\in\mathscr{S}_{0}(\mathcal{O}_{K})$ such that $s_{1}^{\prime}(0)$ is nonzero can have no other fixed points than 0, but many roots. The set of roots of $s_{1}$ and its iterates is denoted by
$\Lambda(s_{1})\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\{x\in\mathfrak{m}_{\mathbb{C}_{p}}\!\mid\exists\,n\geqslant 0,\;s_{1}^{\circ n}(x)=0\}.$
By contrast, a nontorsion invertible power series $s_{2}\in\mathscr{S}_{0}(\mathcal{O}_{K})$ can have no other roots than 0, but many fixed points. When $s_{1}$ and $s_{2}$ commute with each other, Lubin showed in [Lub94, Proposition 3.2] that $\Lambda(s_{1})$ is the set of periodic points of $s_{2}$. Knowing more about these points (their valuation, how many there are for a given valuation) is a key point in our arguments. This can be achieved by means of the Weierstrass preparation theorem and Newton polygons. The use of Newton polygons requires fixing a valuation; for this, we always use the valuation $\upsilon_{K}$ normalized on $K$ so that $\upsilon_{K}(p)=e$.
Proposition 1.1.
For every integer $n\geqslant 1$, the endpoints of the decreasing part of the Newton polygon of $f^{\circ n}\!$ are the $(p^{i},e(n-i))$ for each integer $i\in\{0,\dots,n\}$.
Proof.
We proceed by induction on $n$. The base case ($n=1$) follows directly from the hypotheses on $f$ and [Li96, Theorem 4.1]. The decreasing part of the Newton polygon of $f$ consists of a single line of slope e/(p1)-e/(p-1), implying that the Weierstrass polynomial associated to $f(X)/X$ is irreducible over $K$ since (e,p1)=1(e,p-1)=1. Assume now that the result holds for $n\geqslant 1$. By multiplicativity of the linear coefficients and the Weierstrass degrees,
$(1,e(n+1))\quad\text{and}\quad(p^{n+1},0)$
are vertices of the Newton polygon of $f^{\circ n+1}$. By varying $\pi$ among the $p$ roots of $f$, the equations $f^{\circ n}-\pi=0$ describe all the roots of $f^{\circ n+1}$. When $\pi=0$, one then retrieves the roots of $f^{\circ n}$. Otherwise, the Newton polygon of $f^{\circ n}-\pi$ consists of a single line of slope $e/(p^{n+1}-p^{n})$, which is greater than those of the Newton polygon of $f^{\circ n}$ and hence contributes to the last segment. Packing the roots of $f^{\circ n+1}/f^{\circ n}$, that is, the roots of $f^{\circ n+1}$ that are not roots of $f^{\circ n}$, one thus finds that there are $p^{n}(p-1)$ of them and that the Weierstrass polynomial associated to this segment is irreducible over $K$ since $(e,p^{2}-p)=1$. The left-hand side part of the Newton polygon of $f^{\circ n+1}$ comes from the induction hypothesis ∎
Remark 1.2.
This proof yields two observations:
  1. (1)
    Unlike Specter’s case, the Weierstrass polynomial of $f^{\circ n+1}/f^{\circ n}$ is not Eisenstein in general but still irreducible over $K$, which is sufficient: therefore, the absolute Galois group $G_{K}$ of $K$ acts transitively on the roots of $f^{\circ n+1}/f^{\circ n}$ given that $f$ is defined over $\mathcal{O}_{K}$.
  2. (2)
    Asking for the simplicity of the roots of $f$ and its iterates is equivalent to asking that the p1p-1 nonzero roots of $f$ have the same valuation, under the coprimality assumption. This condition is easier to check in practice because it only requires to draw a finite part of the Newton polygon of $f$.
Corollary 1.3.
For each nonzero $m\in\mathbb{Z}_{p}$, the fixed points of $u^{\circ m}\!$ are the roots of $f^{\circ v(m)}$ where $v(m)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\upsilon_{K}(u^{\prime}(0)^{m}-1)/e$.
Proof.
The series $u$ commutes with the noninvertible $f$ having finite Weierstrass degree, ensuring that all of its nontrivial iterates have finitely many fixed points [Lub94, Corollary 4.3.1]. The decreasing part of the Newton polygon of $u^{\circ m}(X)-X$ evolves between the vertex $(1,\upsilon_{K}(u^{\prime}(0)^{m}-1))$ and the $X$-axis. Since $u$ is defined over $\mathcal{O}_{K}$, its fixed points are a finite union of $G_{K}$-orbits in $\Lambda(f)$. According to Remark 1.2, each segment in the latter polygon then makes a decrease of $e$ along the $Y$-axis. ∎
2. The latent Tate module
As stated in the introduction, the $p$-adic Tate module of a formal group characterizes it up to isomorphism. Specter provided a method to recover the associated representation in [Spe18]. In this section, we adapt this method to our case, and go one step further by retrieving the full structure on the set of $f$-consistent sequences.
2.1. The Galois action on $f$-consistent sequences
The full structure on $T_{f}$ can be recovered solely from its Galois action. To this end, we are interested in finding a more explicit description of this action. This can be achieved by comparing it with the action of the $\mathbb{Z}_{p}$-iterates of $u$ on $T_{f}$. Recall that $T_{f}$ is the $G_{K}$-set of $f$-consistent sequences of elements in $\mathfrak{m}_{\mathbb{C}_{p}}$, and that the pair of series $(f,u)$ is a pair of endomorphisms of it, acting by componentwise evaluation:
$T_{f}=\big\{(x_{n})_{n\geqslant 0}\in\mathfrak{m}_{\mathbb{C}_{p}}^{\mathbb{N}}\!\;\big|\;x_{0}=0\text{ and }\forall\,n\geqslant 0,\;f(x_{n+1})=x_{n}\big\}.$
Let $S_{f}\subset T_{f}$ be the subset composed of all elements whose second entry is nonzero. If Lubin’s conjecture holds, then $T_{f}$ is canonically isomorphic to the $p$-adic Tate module of the latent formal group, and an element $x\in T_{f}$ is a generator of $T_{f}$ as $\mathbb{Z}_{p}$-module if and only if $x\in S_{f}$. Elements of $T_{f}$ are either zero or have finitely many initial zero entries followed by an element of $S_{f}$. In order to describe the $G_{K}$-action on $T_{f}$, it is therefore sufficient to describe it on the subset $S_{f}$. However, recovering this action dynamically via the $\mathbb{Z}_{p}$-iterates of $u$ requires restricting it. Indeed, $u$, and hence its iterates, fix some initial entries of all the elements of $T_{f}$. We let
. . $\ell\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\upsilon_{K}(u^{\prime}(0)-1)/e\geqslant 1$
which is a positive integer since we assume $u^{\prime}(0)\in 1+p\mathbb{Z}_{p}$. According to Corollary 1.3, $u$ fixes up to the $\ell+1$ first components of the elements of $S_{f}$. We view $S_{f}$ as a directed tree:
$\vdots$$\vdots$$\vdots\;$$\vdots$$\vdots$$\vdots\;$$\vdots$$\vdots$$\vdots\;$$\vdots$$\vdots$$\vdots\;$$\vdots$$\vdots$$\vdots\;$$\vdots$$\vdots$$\vdots\;$$\ell$$\ast$$\ast$$\ast$$\ast$$\ast$$\ast$$\ast$$\ast$}0{0}
where the diagram is given in the case $p=3$ and $\ell=2$. Indexed from bottom to top, its vertices at level $n\geqslant 1$ are the roots of $f^{\circ n}/f^{\circ n-1}$, and there is an edge from a vertex $v_{2}$ to a vertex $v_{1}$ if and only if $f(v_{2})=v_{1}$. Given $\pi=(\pi_{n})_{n\geqslant 0}\in S_{f}$, we let
$S_{f,\pi}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\{(x_{n})_{n\geqslant 0}\in T_{f}\mid x_{0}=\pi_{0},\dots,x_{\ell}=\pi_{\ell}\}$
which is a subset of $S_{f}$. We set $j\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=p^{\ell}-p^{\ell-1}$; then any sequence $(\alpha_{1},\dots,\alpha_{j})$ of elements $\alpha_{i}=(\alpha_{i,n})_{n\geqslant 0}$ of $S_{f}$ such that $\{\alpha_{1,\ell},\dots,\alpha_{j,\ell}\}$ is the set of roots of $f^{\circ\ell}/f^{\circ\ell-1}$ determines a partition $(S_{f,\alpha_{1}},\dots,S_{f,\alpha_{j}})$ of $S_{f}$. In what follows, we first retrieve the Galois action on such subsets, and then discuss of its global action on $S_{f}$. The next result allows us to pass from one element of $S_{f}$ to another.
Proposition 2.1.
The Galois group $G_{K}\!$ acts transitively on the set $S_{f}$.
Proof.
For every $n\geqslant 1$, the $n$-th component of an element of $S_{f}$ is a root of $f^{\circ n}$ that is not a root of $f^{\circ n-1}$. These roots are the roots of an irreducible polynomial over $K$, as stated in Remark 1.2. Therefore $G_{K}$ acts transitively on this set and, since $f$ is defined over $\mathcal{O}_{K}$, this action passes to the limit as a transitive action. ∎
2.1.1. Action on cells
Let $\pi=(\pi_{n})_{n\geqslant 0}$ be an element of $S_{f}$, and set $L_{\pi}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=K(\pi_{\ell})$. Then the Galois group $G_{L_{\pi}}\!$ fixes the $\ell+1$ first components of $\pi$. We begin by recovering the action of $G_{L_{\pi}}\!$ on the cell $S_{f,\pi}\subset S_{f}$.
Proposition 2.2.
The group $u^{\circ\mathbb{Z}_{p}}\!$ acts simply transitively on the set $S_{f,\pi}$.
Proof.
This is the same as [Spe18, Proposition 2.9] adapting the notations; the key point lies in Corollary 1.3. For the reader’s convenience and due to the stature of this result, we outline the proof. To show that the action is transitive, it suffices to show that it is the case on
$O_{n}(\pi_{\ell})\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\{x\in\mathfrak{m}_{\mathbb{C}_{p}}\!\mid f^{\circ n-\ell}(x)=\pi_{\ell}\}\subset\Lambda(f)$
for all integers $n>\ell$. Using the Newton polygon, one finds that the cardinality of $O_{n}(\pi_{\ell})$ is $p^{n-\ell}$. Joint to Corollary 1.3, the stabilizer of an element $x\in O_{n}(\pi_{\ell})$ under the action of the $\mathbb{Z}_{p}$-iterates of $u$ is
$\operatorname{Stab}(x)=\big\{u^{\circ m}\in u^{\circ\mathbb{Z}_{p}}\;\big|\;\upsilon_{K}(u^{\prime}(0)^{m}-1)\geqslant ne\big\}=u^{\circ p^{n-\ell}\mathbb{Z}_{p}}$
since we have picked $u^{\prime}(0)\in 1+p\mathbb{Z}_{p}$. By the orbit-stabilizer theorem, we obtain that $u^{\circ\mathbb{Z}_{p}}$ acts transitively on $O_{n}(\pi_{\ell})$. Consequently, the action of the statement is also free. ∎
The transitive action of $G_{K}$ on the set $S_{f}$ restricts to a transitive action of $G_{L_{\pi}}\!$ on the subset $S_{f,\pi}$. Combining this observation with Proposition 2.2 enables us to recover the latter action via the following rule: as $g$ varies in $G_{L_{\pi}}$, the elements $g.\pi$ exhaust the set $S_{f,\pi}$, and comparing $g.\pi$ with $\pi$ gives a unique $u_{g}\in u^{\circ\mathbb{Z}_{p}}$ satisfying
$g.\pi=u_{g}(\pi).$
The resulting map sending $g$ to $u_{g}^{\prime}(0)$ defines a character.
Proposition 2.3.
Using the previous notation, the map $\chi_{f,\pi}\colon G_{L_{\pi}\!}\rightarrow\mathbb{Z}_{p}^{\times}\colon g\mapsto u_{g}^{\prime}(0)$ is a character of $G_{L_{\pi}}\!$ satisfying $g.\pi=[\chi_{f,\pi}(g)]_{f}(\pi)$ for every $g\in G_{L_{\pi}}$.
Proof.
The fact that $\chi_{f,\pi}$ is a group homomorphism is a consequence of the freeness of the action in Proposition 2.2. The rest of the statement is immediate. ∎
2.1.2. Action on the whole set
Having analyzed the situation on cells, we turn to the global case. To compare the local actions with each other, we consider the finite extension $L$ of $K$ generated by the roots of $f^{\circ\ell}/f^{\circ\ell-1}$, or equivalently, the Galois closure of $L_{\pi}/K$ for any $\pi\in S_{f}$. Given $\pi,\pi^{\prime}\in S_{f}$, Proposition 2.1 guarantees the existence of a $\sigma\in G_{K}$ satisfying $\pi^{\prime}=\sigma.\pi$. Consequently $L_{\pi^{\prime}}=\sigma(L_{\pi})$ and the absolute Galois groups of $L_{\pi}$ and $L_{\pi^{\prime}}$ are $\sigma$-conjugate in $G_{K}$. Since $u$ is defined over $\mathcal{O}_{K}$, the previous rule gives, for every $g\in G_{L}$, that
(2) $\chi_{f,\pi}(g)=\chi_{f,\pi^{\prime}}(\sigma g\sigma^{-1}).$
Now if Lubin’s conjecture holds, we should also obtain equality of the characters, since there would exist a unique character giving the action of $G_{K}$ at every point of $T_{f}$. This lack of precision in our situation reflects the price we pay for restricting the Galois action from $K$ to $L$. Nevertheless, we fix once and for all a $\pi=(\pi_{n})_{n\geqslant 0}\in S_{f}$ and let
$\chi_{f}\colon G_{L}\longrightarrow\mathbb{Z}_{p}^{\times}\colon g\longmapsto\chi_{f,\pi}(g).$
After the two preparatory subsections that follow, we shall prove that this character is crystalline of weight $1$. Therefore, there exists a formal group over $\mathcal{O}_{L}$ whose associated character is $\chi_{f}$, and we transport the structure of its Tate module to $T_{f}$. Theorem 3.1 will ultimately imply that the choice of $\pi$ was independent of the subsequent constructions.
2.2. Fontaine’s period rings
Fontaine achieved a classification of the $p$-adic representations of the absolute Galois group $G_{E}$ of a finite extension $E/\mathbb{Q}_{p}$. His classifier relies on rings such as $B_{\mathrm{dR}}$ and $B_{\mathrm{cris}}$. In this subsection, we outline the construction of these two rings and review the formal group setting. The original constructions can be found in [Fon94] and in [Col02] for the tensorial generalization. To coincide with our situation, we work directly in the particular case of $E=L$.
2.2.1. The integers of the tilt of $\mathbb{C}_{p}$
Let $\mathcal{O}_{\mathbb{C}_{p}}\!$ be the closed unit disk of $\mathbb{C}_{p}$. Throughout the remainder of this article, we fix $\omega_{L}$ a uniformizer of $L$ and let $q$ be the cardinality of the residue field $\mathcal{O}_{L}/\mathfrak{m}_{L}=\mathbb{F}_{q}$. The definition of the aforementioned period rings involves several mechanisms, whose starting point is the perfect ring
[…]
of characteristic $p$. It is endowed with a valuation $\upsilon_{\mathcal{R}}$ for which it is complete, defined by $\upsilon_{\mathcal{R}}(x)=\lim_{n\to+\infty}q^{n}\upsilon_{p}(\widehat{x}_{n})$ for all $x=(x_{n})_{n\geqslant 0}\in\mathcal{R}$, where $\widehat{x}_{n}\in\mathcal{O}_{\mathbb{C}_{p}}\!$ is a lift of $x_{n}$ modulo $(\omega_{L})$ for all integers $n\geqslant 0$ and $\upsilon_{p}$ is the $p$-adic valuation normalized on $\mathbb{Q}_{p}$. This ring is also equipped with a compatible action of $G_{L}$ inherited from its action on $\mathcal{O}_{\mathbb{C}_{p}\!}$.
2.2.2. Blank period rings
We now pass to characteristic zero. Let $\operatorname{W}_{\!\omega_{L}\!}$ be the generalized $\omega_{L}$-adic Witt vectors endofunctor of $\mathcal{O}_{L}$-algebras; see [CD15, §2] for a presentation. The ring $\mathcal{R}$ is naturally an $\mathcal{O}_{L}$-algebra and we let
$A_{\mathrm{inf},L}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{W}_{\!\omega_{L}\!}(\mathcal{R})$
be the unique strict $\omega_{L}$-ring of residue ring $\mathcal{R}$. Alternatively, one can see this blank period ring as $\mathcal{O}_{L}\otimes_{\operatorname{W}(\mathbb{F}_{q})}\operatorname{W}(\mathcal{R})$, where $\operatorname{W}$ is the usual $p$-adic Witt vectors functor. Elements of this ring can be uniquely written in the form $\sum_{n\geqslant 0}[x_{n}]\omega_{L}^{n}$ where $x_{n}\in\mathcal{R}$ and [][-] is the Teichmüller lift. The action of $G_{L}$ on $\mathcal{R}$ lifts by functoriality to an action on $A_{\mathrm{inf},L}$. An important map attached to it is the $G_{L}$-equivariant epimorphism $\theta\colon A_{\mathrm{inf},L}\twoheadrightarrow\mathcal{O}_{\mathbb{C}_{p}}\!$ of $\mathcal{O}_{L}$-algebras uniquely determined by mapping $[x]$ to $\lim_{n\to+\infty}\widehat{x}_{n}^{q^{n}}\!$, where $\widehat{x}_{n}\in\mathcal{O}_{\mathbb{C}_{p}}\!$ is a lift of the $n$-th component of $x\in\mathcal{R}$. This map extends in a unique way by $L$-linearity to a $G_{L}$-equivariant epimorphism $\theta$ from $B_{\mathrm{inf},L}^{+}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=A_{\mathrm{inf},L}[1/\omega_{L}]$ to $\mathbb{C}_{p}$, whose kernel is principal.
2.2.3. The de Rham period field
The de Rham period ring
[…]
is the $\ker(\theta)$-adic completion of $B_{\mathrm{inf},L}^{+}$. It inherits the topology generated by the intersection of the $\ker(\theta)$-adic and $\omega_{L}$-adic topologies from its dense subring of definition, along with a compatible action of $G_{L}$. Once again, $\theta$ extends to a $G_{L}$-equivariant epimorphism $\theta\colon B_{\mathrm{dR}}^{+}\twoheadrightarrow\mathbb{C}_{p}$ and its kernel is still principal. The $\mathbb{Z}$-powers of $\ker(\theta)$ induce a decreasing, exhaustive and separated filtration $(\operatorname{Fil}^{n}B_{\mathrm{dR}})_{n\in\mathbb{Z}}$ on the fraction field $B_{\mathrm{dR}}$ of $B_{\mathrm{dR}}^{+}$. This field is our first ring of interest.
The field $B_{\mathrm{dR}}$ is sufficient to capture $p$-adic representations arising from formal groups, but too coarse to be used to classify them. Denote by $\operatorname{\mathbf{Rep}}_{\mathbb{Q}_{p}}(G_{L})$ the category of $p$-adic representations of $G_{L}$, and by $\operatorname{\mathbf{Fil}}_{L}$ the one of finite-dimensional $L$-vector spaces equipped with a decreasing, exhaustive and separated $\mathbb{Z}$-filtration by subspaces. Fontaine defined a contravariant functor
$\operatorname{\mathbf{D}}_{\mathrm{dR}}^{*}\colon\operatorname{\mathbf{Rep}}_{\mathbb{Q}_{p}}(G_{L})\longrightarrow\operatorname{\mathbf{Fil}}_{L}$
by sending a $p$-adic representation $V$ to $\operatorname{\mathbf{D}}_{\mathrm{dR}}^{*}(V)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{Hom}_{\mathbb{Q}_{p}[G_{L}]}(V,B_{\mathrm{dR}})$, equipped with the induced filtration $\operatorname{Fil}^{n}\operatorname{\mathbf{D}}_{\mathrm{dR}}^{*}(V)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{Hom}_{\mathbb{Q}_{p}[G_{L}]}(V,\operatorname{Fil}^{n}B_{\mathrm{dR}})$ for all $n\in\mathbb{Z}$. We say that a $p$-adic representation $V$ of $G_{L}$ is de Rham if the $L$-dimension of $\operatorname{\mathbf{D}}_{\mathrm{dR}}^{*}(V)$ is equal to the $\mathbb{Q}_{p}$-dimension of $V$. In this case, the jumps of the filtration are called the Hodge-Tate weights of $V$. Restricting the preceding functor to de Rham representations yields a faithful functor. When $(V,\rho)$ is one-dimensional, being de Rham is equivalent to the existence of a nonzero element $t\in B_{\mathrm{dR}}$ such that $g.t=\rho(g)t$ for all $g\in G_{L}$. Then, the Hodge-Tate weight is the largest $w\in\mathbb{Z}$ such that $t\in\operatorname{Fil}^{w}B_{\mathrm{dR}}$.
2.2.4. The crystalline period ring
To refine the theory, Fontaine defined a subring $B_{\mathrm{cris}}$ of $B_{\mathrm{dR}}$ as follows: let $A_{\mathrm{cris}}^{\circ}$ be the PD-envelope of $\operatorname{W}(\mathcal{R})$ with respect to $\ker(\theta)$. Set $A_{\mathrm{cris}}$ to be its $p$-adic completion. One can show that the latter is still a subring of $B_{\mathrm{dR}}$. Define $B_{\mathrm{cris}}^{+}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=A_{\mathrm{cris}}[1/p]$. The $G_{L}$-action on $\operatorname{W}(\mathcal{R})$ extends at each previous level and the latter contains a unique $G_{L}$-stable $\mathbb{Q}_{p}$-line on which $G_{L}$ acts by the cyclotomic character. The crystalline period ring
$B_{\mathrm{cris}}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=B_{\mathrm{cris}}^{+}[1/t]$
is obtained from $B_{\mathrm{cris}}^{+}$ by inverting any nonzero $t$ in this line. Unlike $B_{\mathrm{dR}}$, the Frobenius $\varphi$ on $\operatorname{W}(\mathcal{R})$ extends uniquely to $B_{\mathrm{cris}}$ by setting $\varphi(t)=pt$, thereby yielding an injective $G_{L}$-equivariant endomorphism $\varphi\colon B_{\mathrm{cris}}\hookrightarrow B_{\mathrm{cris}}$ which is semilinear with respect to the Frobenius $\sigma$ of $L_{0}$, the maximal unramified extension of $\mathbb{Q}_{p}$ inside $L$. The canonical map $L\otimes_{L_{0}}B_{\mathrm{cris}}\rightarrow B_{\mathrm{dR}}$ is injective.
The ring $B_{\mathrm{cris}}$ will be of primary interest to us, as it is rich enough not only to capture but also to classify representations arising from formal groups. Let $\operatorname{\mathbf{MF}}_{L}(\varphi)$ denote the category of finite-dimensional $L_{0}$-vector spaces equipped with a bijective $\sigma$-semilinear Frobenius endomorphism $\varphi$, whose scalar extension to $L$ is an object of $\operatorname{\mathbf{Fil}}_{L}$. Morphisms in this category are $L_{0}$-linear maps commuting with $\varphi$ and preserving the filtration after scalar extension. Fontaine defined a contravariant functor
$\operatorname{\mathbf{D}}_{\mathrm{cris}}^{*}\colon\operatorname{\mathbf{Rep}}_{\mathbb{Q}_{p}}(G_{L})\longrightarrow\operatorname{\mathbf{MF}}_{L}(\varphi)$
by sending a $p$-adic representation $V$ to $\operatorname{\mathbf{D}}_{\mathrm{cris}}^{*}(V)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{Hom}_{\mathbb{Q}_{p}[G_{L}]}(V,B_{\mathrm{cris}})$, equipped with the Frobenius induced by that on $B_{\mathrm{cris}}$. Moreover, one has $L\otimes_{L_{0}}\operatorname{\mathbf{D}}_{\mathrm{cris}}^{*}(V)\simeq\operatorname{\mathbf{D}}_{\mathrm{dR}}^{*}(V)$. We say that a $p$-adic representation $V$ of $G_{L}$ is crystalline if the $L_{0}$-dimension of $\operatorname{\mathbf{D}}_{\mathrm{cris}}^{*}(V)$ is equal to the $\mathbb{Q}_{p}$-dimension of $V$. In this case, its Hodge-Tate weights are its weights as a de Rham representation. In this setting, the previous functor yields an antiequivalence of categories
$\operatorname{\mathbf{D}}_{\mathrm{cris}}^{*}\colon\operatorname{\mathbf{Rep}}_{\mathbb{Q}_{p}}^{\mathrm{cris}}(G_{L})\longrightarrow\operatorname{\mathbf{MF}}_{L}^{\mathrm{adm}}(\varphi)$
where the superscript “cris” stands for the crystalline ones and “adm” for the admissible ones. Concretely, a one-dimensional filtered $\varphi$-module $D$ with jump at $j\in\mathbb{Z}$ is admissible if and only if $j=\upsilon_{p}(\varphi(d)/d)$ where $d\in D$ is any nonzero vector.
2.2.5. Application to formal groups
Given a formal group $H$ of height $h\geqslant 1$ over $\mathcal{O}_{L}$, the $\mathbb{Q}_{p}$-scalar extension of its $p$-adic Tate module
$\operatorname{V}_{\!p}(H)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}\operatorname{T}_{p}(H)$
is a de Rham $G_{L}$-representation of dimension $h$ and of Hodge-Tate weight $1$. It is even a crystalline representation of $G_{L}$. In this paper, we are primarily interested in the integral setting and the situation is as follows. Let $\operatorname{\mathbf{FGL}\!}/\mathcal{O}_{L}$ denote the category of formal groups of finite height defined over $\mathcal{O}_{L}$, and $\operatorname{\mathbf{Rep}}_{\mathbb{Z}_{p}}^{\mathrm{cris},1}(G_{L})$ the category of $G_{L}$-stable $\mathbb{Z}_{p}$-lattices in crystalline representations of $G_{L}$ of Hodge-Tate weights $1$. By [SW13, Corollary 6.2.3] and [Tat67, Proposition 1], the functor
$\operatorname{T}_{p}\colon\operatorname{\mathbf{FGL}\!}/\mathcal{O}_{L}\longrightarrow\operatorname{\mathbf{Rep}}_{\mathbb{Z}_{p}}^{\mathrm{cris},1}(G_{L})$
is an equivalence of categories. When $H$ is of height $1$, any endomorphism $s\in\operatorname{End}_{\mathcal{O}_{L}}(H)$ acts on $\operatorname{T}_{p}(H)$ by multiplication by its linear coefficient $s^{\prime}(0)\in\mathbb{Z}_{p}$, and hence extends to an endomorphism of $\operatorname{V}_{\!p}(H)$. In view of the equivalence and the fact that this scalar extension is faithful, the single coefficient $s^{\prime}(0)$ together with the $\mathbb{Z}_{p}$-module structure on $\operatorname{T}_{p}(H)$ or $\operatorname{V}_{\!p}(H)$ uniquely determine the endomorphism $s$. It also acts on $\operatorname{\mathbf{D}}_{\mathrm{dR}}^{*}(\operatorname{V}_{\!p}(H))$ in the following way: for a nonzero $v\in\operatorname{V}_{\!p}(H)$ and a nonzero period $t\in B_{\mathrm{dR}}$ of $\operatorname{V}_{\!p}(H)$, the map induced by $s$ sends $(v\mapsto t)$ to $(v\mapsto s^{\prime}(0)t)$ since $s(v)=s^{\prime}(0)v$. The initial series $s$ is unequivocally determined by this action.
2.3. The universal $f^{\circ r}$-consistent ring
Denote by $r$ the residue degree of $L/\mathbb{Q}_{p}$, so that $q=p^{r}$. To suit this parameter, we adapt the universal $f$-consistent ring defined in [Spe18] by replacing $f$ with its $r$-th iterate. According to [Lub94, Corollary 6.2.1], there exists an invertible power series $a\in\mathscr{S}_{0}(\mathbb{F}_{q})$ satisfying
$f^{\circ r}(X)\bmod(\omega_{L})=a(X^{q}).$
We consider the colimit
[…]
with respect to the identification $X_{n}\mapsto f^{\circ r}(X_{n+1})$ for all integers $n\geqslant 0$, and we denote by $A_{\infty}(\mathcal{O}_{L})$ its $\omega_{L}$-adic completion. We endow this ring with an action of $G_{L}$ by letting $g.X_{n}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=[\chi_{f}(g)]_{f}(X_{n})$ for every integer $n\geqslant 0$.
Proposition 2.4.
The ring $A_{\infty}(\mathcal{O}_{L})$ is a strict $\omega_{L}$-ring.
Proof.
By construction, the ring $A_{\infty}(\mathcal{O}_{L})$ is $\omega_{L}$-adically complete, and $\omega_{L}$ is not a zero divisor. To show that $A_{\infty}(\mathcal{O}_{L})/(\omega_{L})$ is a perfect ring of characteristic $p$, observe that
[…]
Universal property.
Endowed with the $(\omega_{L},X_{0})$-adic topology, every formal variable $X_{n}$ is topologically nilpotent in $A_{\infty}(\mathcal{O}_{L})$. Moreover, this ring is universal for $f^{\circ r}$-consistent sequences in the following sense: given a pair $(R,s)$, where $R$ is an $\omega_{L}$-adically complete, ind-complete adic $\mathcal{O}_{L}$-algebra and $s=(s_{n})_{n\geqslant 0}$ is a $f^{\circ r}$-consistent sequence of topologically nilpotent elements in $R$, there exists a unique homomorphism $\phi_{s}\colon A_{\infty}(\mathcal{O}_{L})\rightarrow R$ sending each $X_{n}$ to $s_{n}$ for every $n\geqslant 0$. Conversely, any homomorphism from $A_{\infty}(\mathcal{O}_{L})$ to $R$ arises in this way for some sequence $s$.
2.4. Regularity of the Galois action
We prove that the character $\chi_{f}=\chi_{f,\pi}$ chosen in §2.1 is crystalline of Hodge-Tate weight $1$. For this, we pass to the ring $\mathcal{R}$ via the map
$\vartheta_{f}\colon T_{f}\longrightarrow\mathcal{R}\colon(x_{n})_{n\geqslant 0}\longmapsto(a^{\circ n}(x_{nr}\bmod\omega_{L}))_{n\geqslant 0}.$
Then we apply the universal mapping property of the ring $A_{\infty}(\mathcal{O}_{L})$ to the pair $(\mathcal{R},\varpi)$, where $\varpi=(\varpi_{n})_{n\geqslant 0}$ is the $f^{\circ r}$-consistent sequence of topologically nilpotent elements in $\mathcal{R}$ given by $\varpi_{n}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=(a^{\circ m}(\pi_{(n+m)r}\bmod\omega_{L}))_{m\geqslant 0}$ for each $n\geqslant 0$. Let $\phi_{\varpi}\colon A_{\infty}(\mathcal{O}_{L})\rightarrow\mathcal{R}$ be the corresponding homomorphism. As $\omega_{L}$ is zero in $\mathcal{R}$, this map factors through $(\omega_{L})$; it is even its kernel.
Proposition 2.5.
The homomorphism $\phi_{\varpi}$ is $G_{L}$-equivariant with kernel $\omega_{L}A_{\infty}(\mathcal{O}_{L})$.
Proof.
The first assertion follows from the fact that all the power series $\chi_{f}(-)]_{f}$ commute with $f$. The second is the same as in [Ber16, Lemma 2.2]: if the induced map $\mathbb{F}_{q}[\![X_{n}]\!]\rightarrow\mathcal{R}$ were not injective, there would exist a nonzero $P(X_{n})\in\mathbb{F}_{q}[X_{n}]$ such that $P(\overline{\phi_{\varpi}}(X_{n}))=0$ and this would imply that $\varpi_{n}$ is algebraic over $\mathbb{F}_{p}$, which is not the case. ∎
Proposition 2.4 asserts that $A_{\infty}(\mathcal{O}_{L})$ is a strict $\omega_{L}$-ring, hence the factorization of $\phi_{\varpi}$ through its kernel lifts to an injective $G_{L}$-equivariant homomorphism $A_{\infty}(\mathcal{O}_{L})\hookrightarrow A_{\mathrm{inf},L}$. Under this embedding, we let $x_{n}$ be the image of $X_{n}$ for all integers $n\geqslant 0$. The action of the group $G_{L}$ on these points is particularly interesting, as we shall prove that applying the $f$-logarithm to them produces periods of $\chi_{f}$. As it stands, by [Ber14, Lemma 3.2], we already know that
$t_{f}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{Log}_{f}(x_{0})\in B_{\mathrm{dR}}^{+}.$
This series converges in $B_{\mathrm{dR}}^{+}$ because $\operatorname{Log}_{f}$ converges on $\mathfrak{m}_{\mathbb{C}_{p}}\!$ and $\theta(x_{n})\equiv\pi_{nr}\bmod(\omega_{L})$ for all $n\geqslant 0$. Considering $\operatorname{Log}_{f}(x_{n})$ for a different integer $n>0$ would yield a $\mathbb{Q}_{p}$-multiple of $t_{f}$. Note that $t_{f}$ depends not only on $f$ but also on the choice of the consistent sequence $\pi$ in subsection 2.1. We proceed to show that it is a nonzero period of $\chi_{f}$.
Proposition 2.6.
The character $\chi_{f}\!$ of $G_{L}\!$ is crystalline of Hodge-Tate weight $1$.
Proof.
The previous convergence in $B_{\mathrm{dR}}^{+}$ is with respect to the $\ker(\theta)$-adic topology. So there exists some $m\geqslant 1$ for which $\theta(b_{m}x_{0}^{m})=b_{m}\theta(x_{0})^{m}=0$, implying $\theta(x_{0})=0$, where $b_{m}\in K^{\times}$ is the $m$-th coefficient of the formal power series $\operatorname{Log}_{f}$. Consequently, $\theta(t_{f})=0$ and so $\chi_{f}$ has weight $\geqslant 1$. If it had weight $2$, then the $\upsilon_{\mathcal{R}}$-valuation of $x_{0}\bmod(\omega_{L})=\varpi_{0}$ would be at least $2$, however
$\upsilon_{\mathcal{R}}(\varpi_{0})$ $=\lim_{m\to+\infty}q^{m}\upsilon_{p}(\pi_{mr})$
$=\lim_{m\to+\infty}p^{mr}\frac{1}{p^{mr-1}(p-1)}$
$=\frac{p}{p-1}$
is less than $2$ when $p$ is odd, a contradiction. For $p=2$, it suffices to compute the next component of $x_{0}$ and this leads to a contradiction too. In particular $t_{f}$ is nonzero. For every $g\in G_{L}$, we have
$g.t_{f}=\operatorname{Log}_{f}(g.x_{0})=\operatorname{Log}_{f}([\chi_{f}(g)]_{f}(x_{0}))=\chi_{f}(g)t_{f}$
and hence $\chi_{f}$ is de Rham of Hodge-Tate weight $1$. To show that $\chi_{f}$ is indeed crystalline, it suffices to observe that $t_{f}$ lies in $L\otimes_{L_{0}}B_{\mathrm{cris}}^{+}$ because $\theta(x_{0})=0$ and $\operatorname{Log}_{f}\in\mathscr{S}_{0}(K)$, and its existence implies the existence of a crystalline period since
$\dim_{\mathbb{Q}_{p}}\mathbb{Q}_{p}(\chi_{f})$ $=\dim_{L}\operatorname{Hom}_{\mathbb{Q}_{p}[G_{L}]}(\mathbb{Q}_{p}(\chi_{f}),L\otimes_{L_{0}}B_{\mathrm{cris}})$
$=\dim_{L}L\otimes_{L_{0}}\operatorname{Hom}_{\mathbb{Q}_{p}[G_{L}]}(\mathbb{Q}_{p}(\chi_{f}),B_{\mathrm{cris}})$
$=\dim_{L_{0}}\operatorname{\mathbf{D}}^{*}_{\mathrm{cris}}(\mathbb{Q}_{p}(\chi_{f})).\qed$
Remark 2.7.
While in [Spe18] the period $t_{f}$ is already in $B_{\mathrm{cris}}^{+}$, this is not the case here in general. A major obstruction arises from the fact that the formal power series $\operatorname{Log}_{f}$ is not, a priori, defined over the unramified part of $K/\mathbb{Q}_{p}$ when $e>1$.
2.5. Defining the Tate module structure
Let us summarize the situation up to this point. Starting from $(f,u)$, we have, over a finite extension $L/K$, recovered a restriction of the action of $G_{K}$ on cells of $T_{f}$ via the one of the $\mathbb{Z}_{p}$-iterates of $u$. The chosen resulting character $\chi_{f}$ of $G_{L}$ is such that $\mathbb{Q}_{p}(\chi_{f})$ is crystalline of Hodge-Tate weight $1$, and thus encodes the data of a formal group. Nevertheless, this construction severs the direct link with the initial set $T_{f}$. Our final goal for this section is to relate the information carried by $\mathbb{Z}_{p}(\chi_{f})$ back to the $G_{L}$-set $T_{f}$, thereby endowing it with a $\mathbb{Z}_{p}$-module structure that enables the recovery of the latent formal group.
2.5.1. A pivotal formal group
Since $\chi_{f}$ is crystalline of Hodge-Tate weight $1$, there exists a height-$1$ formal group $H$ over $\mathcal{O}_{L}$ such that $\operatorname{T}_{p}(H)\simeq\mathbb{Z}_{p}(\chi_{f})$, as stated in §2.2.5. This formal group may not be the one we are looking for, but if Lubin’s conjecture holds true, it would be isomorphic to it. We will transport the structure of $\operatorname{T}_{p}(H)$ onto $T_{f}$. Notice that any other choice of formal group $H^{\prime}$ such that $\operatorname{T}_{p}(H^{\prime})\simeq\mathbb{Z}_{p}(\chi_{f})$ would induce the same structure on $T_{f}$.
Lemma 2.8.
The action of $G_{L}\!$ on $\operatorname{T}_{p}(H)$ is given by $\chi_{f}$.
Proof.
Let $\chi$ be the character of $\operatorname{T}_{p}(H)$. Since there exists a $\mathbb{Z}_{p}[G_{L}]$-module isomorphism $\tau\colon\!\operatorname{T}_{p}(H)\rightarrow\mathbb{Z}_{p}(\chi_{f})$, for every $x\in\operatorname{T}_{p}(H)$ and all $g\in G_{L}$, one has $\chi(g)\tau(x)=\chi_{f}(g)\tau(x)$. This implies that $\chi=\chi_{f}$ given that $\mathbb{Z}_{p}(\chi_{f})$ is free over $\mathbb{Z}_{p}$. ∎
We certainly wish $(f,u)$ to remain a pair of endomorphisms of $T_{f}$ after transporting the structure. To this end we identify the unique endomorphisms of $H$ to which $f$ and $u$ should correspond by examining how they act on the constructed period $t_{f}$, which is a shared data. Indeed, $t_{f}$ was constructed from the dynamical system and by Lemma 2.8 it is also a period of $\operatorname{T}_{p}(H)$. Let $s\in\{f,u\}$; therefore $s$ acts naturally on $T_{f}$. We want to define $s(t_{f})$ arising from this action in a way compatible with the construction of $t_{f}$. Recall that in §2.4 we have followed this path in order to obtain a period:
$A_{\infty}(\mathcal{O}_{L})$$A_{\mathrm{inf},L}$$L\otimes_{L_{0}}B_{\mathrm{cris}}^{+}$$X_{0}$$x_{0}$$t_{f}$$T_{f}$$\mathcal{R}$$\pi$$\varpi_{0}$$\scriptstyle{\operatorname{Log}_{f}}$$\scriptstyle{\phi_{\varpi}}$$\scriptstyle{\vartheta_{f}}$
where the dashed line is defined for all $x\in A_{\mathrm{inf},L}\cap\ker(\theta)$. Considering maps in the same order as we did previously, $s(\pi)$ is sent to the element $(a^{\circ m}(s(\pi_{mr})\bmod\omega_{L}))_{m\geqslant 0}$ which is well defined in $\mathcal{R}$ because $s\bmod(\omega_{L})$ commutes with $a$. This element is nothing but $s(\varpi_{0})$ and is the image of $s(X_{0})$ under $\phi_{\varpi}$. Hence these maps agree after action of $s$. By functoriality of the Witt vectors functor, $s(X_{0})$ is sent to $s(x_{0})$ and lifts $s(\varpi_{0})$ modulo $(\omega_{L})$. As $s(x_{0})\in\ker(\theta)$, we let
$s(t_{f})\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{Log}_{f}(s(x_{0}))=s^{\prime}(0)t_{f}.$
This definition is natural and coincides with the action of a formal group endomorphism on its period, as discussed in paragraph 2.2.5. It thus establishes how $f$ and $u$ act on $t_{f}$. From the perspective of $H$, we assert that $s^{\prime}(0)t_{f}$ arises from $s_{H}$, the multiplication-by-$s^{\prime}(0)$ endomorphism of $H$. Indeed, the composition of the functors considered in §2.2 is faithful:
$\operatorname{\mathbf{Rep}}_{\mathbb{Q}_{p}}^{\mathrm{cris},1}(G_{L})$$\operatorname{\mathbf{MF}}_{L}^{\mathrm{adm},1}(\varphi)$$\operatorname{\mathbf{FGL}\!}/\mathcal{O}_{L}$$\operatorname{\mathbf{Rep}}_{\mathbb{Z}_{p}}^{\mathrm{cris},1}(G_{L})$$\scriptstyle{\operatorname{\mathbf{D}}^{*}_{\mathrm{cris}}}$$\scriptstyle{\operatorname{T}_{p}}$$\scriptstyle{\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}-}$
and $s_{H}$ acts in the same way on all de Rham periods of $\operatorname{V}_{\!p}(H)$, including the crystalline ones. We must therefore associate $f$ with $f_{H}$ when constructing a bijection from $T_{f}$ to $\operatorname{T}_{p}(H)$. We also want this bijection to be $G_{L}$-equivariant. However, Lemma 2.8 implies that it is the $\mathbb{Z}_{p}$-iterates of $u_{H}$ that give the global action of $G_{L}$ on $\operatorname{T}_{p}(H)$. This requires us to take into account the weak knowledge (2) in order to preserve the $G_{L}$-action on $T_{f}$, although it does not guarantee that $u$ is associated to $u_{H}$ and hence that it becomes an endomorphism of $T_{f}$ after transport of structure. Nevertheless, it suffices that $f$ be a formal group endomorphism for $u$ to be one as well.
2.5.2. Transport of structure
Given that $H$ has height $1$, its group of points in $\mathfrak{m}_{\mathbb{C}_{p}}\!$ has exactly $p^{n}$ points of $p^{n}$-torsion, which are precisely the roots of $f^{\circ n}_{H}$, for all integers $n\geqslant 0$. To facilitate the construction of a bijection from $T_{f}$ to $\operatorname{T}_{p}(H)$, we reinterpret the latter Tate module by taking the transition maps to be $f_{H}$ instead of $p]_{H}$, which is legitimate because $f^{\prime}(0)$ is a uniformizer in $\mathbb{Z}_{p}$. Considering the above facts, it is clear that such a bijection exists. We now proceed to construct one satisfying our constraints.
Proposition 2.9.
There exists a $\mathbb{Z}_{p}[G_{L}]$-module structure on the set $T_{f}\!$ for which it is a crystalline character of $G_{L}\!$ of Hodge-Tate weight $1$, and $f\in\operatorname{End}_{\mathbb{Z}_{p}[G_{L}]}(T_{f})$.
Proof.
Analogously to the constructions in §2.1, we define $S_{H}\subset\operatorname{T}_{p}(H)$ to be the subset of elements with a nonzero second entry. Given $\beta=(\beta_{n})_{n\geqslant 0}\in S_{H}$, we let
[…]
In order to define a $G_{L}$-equivariant bijection $\tau\colon T_{f}\rightarrow\operatorname{T}_{p}(H)$ transporting $f$ to $f_{H}$, we begin by matching $S_{f}$ with $S_{H}$. Set $j\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=p^{\ell}-p^{\ell-1}$ and choose a system $(\alpha_{1},\dots,\alpha_{j})$ of elements of $S_{f}$ inducing a partition of $S_{f}$ as in §2.1. Similarly, pick a system $(\beta_{1},\dots,\beta_{j})$ of elements of $S_{H}$ where we run through the torsion points of $H$ of exact order $p^{\ell}$. Such a system, in turn, induces a partition of $S_{H}$. For each integers $1\leqslant i\leqslant j$ and $n\geqslant 0$, and all $g\in G_{L}$, set
(3) $\tau((f^{\circ n}\circ[\chi_{f,\alpha_{i}}(g)]_{f})(\alpha_{i}))\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=(f_{H}^{\circ n}\circ[\chi_{f}(g)]_{H})(\beta_{i})$
and $\tau(0)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=0$. Fixing $n=0$ and taking $i\in\{1,\dots,j\}$ in (3) gives a well-defined bijection $\tau\colon S_{f,\alpha_{i}}\rightarrow S_{H,\beta_{i}}$ since the $\mathbb{Z}_{p}$-iterates of $u$ act simply transitively on $S_{f,\alpha_{i}}$, and the same holds with $u_{H}$ and $S_{f,\beta_{i}}$. Varying $n\geqslant 0$ in (3) allows us to reach the other points of $T_{f}$ and $\operatorname{T}_{p}(H)$. This defines a $G_{L}$-equivariant bijection $\tau\colon T_{f}\rightarrow\operatorname{T}_{p}(H)$ associating $f$ with $f_{H}$, and the statement follows by transport of structure. ∎
Remark 2.10.
A different choice of representatives for the partitions in the previous proof would yield the same $\mathbb{Z}_{p}[G_{L}]$-module structure on $T_{f}$, but would permute the roles of the individual elements. This is immaterial for our purpose, since what matters to us are the group law and the property that the power series $f$ remains an endomorphism of $T_{f}$ after transport of structure, which are unaffected by such choices. Nevertheless, the latent formal group $F$ obtained in Theorem 3.1 will satisfy $\operatorname{T}_{p}(F)=T_{f}$ as $G_{L}$-sets, but only $\operatorname{T}_{p}(F)\simeq T_{f}$ as $\mathbb{Z}_{p}[G_{L}]$-modules.
3. The latent formal group
Possessing a $p$-adic Tate module and having access to the existence of a quasi-inverse to the equivalence in §2.2.5 is not sufficient to solve Lubin’s conjecture: it is not enough to work up to isomorphism; one must have equality. For this purpose, each time a functor is applied to $T_{f}$, it is necessary to keep track of the behavior of $f$. Recall that $f$ acts on $T_{f}$ by componentwise evaluation, or equivalently, by multiplication by its linear coefficient, and that all its other coefficients are uniquely determined by the action of $\mathbb{Z}_{p}$ on $T_{f}$. Our strategy requires a careful analysis of the explicit constructions in integral $p$-adic Hodge theory. One such approach proceeds through the work of Kisin in [Kis06] and [Kis09] to obtain a formal Breuil-Kisin module, and then via the one of Zink in [Zin01] and [Zin02] on displays, indicated by the dashed path in the diagram:
[…]
where the categories and functors appearing therein will be defined progressively throughout this section. At each stage, we track the action of $f$ on the object being constructed and ensure that it indeed encodes the data of its other coefficients just as $T_{f}$ does. To this end, we construct inside or extend the previous object. These precautions are necessary in order to conclude that the formal group obtained in the end is the latent one.
3.1. The associated Breuil-Kisin module
Recall that we denote by $L_{0}$ the maximal unramified extension of $\mathbb{Q}_{p}$ inside $L$, and $\sigma$ its Frobenius. We have also fixed a uniformizer $\omega_{L}$ of $L$, and we let $E(X)$ denote its Eisenstein minimal polynomial over $L_{0}$. Let $L_{\infty}\!$ be the extension generated over $L$ by the components of a compatible sequence of $p$-th roots of $\omega_{L}$ in $\mathbb{C}_{p}$. We define two subrings $\mathfrak{S}\subset\mathfrak{S}^{\mathrm{ur}}$ of $\operatorname{W}(\mathcal{R})$ as follows. Let
$\mathfrak{S}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{W}(\mathbb{F}_{q})[\![X]\!]$
be the Kisin ring, endowed with a $\sigma$-semilinear Frobenius endomorphism $\varphi$ defined by $\varphi(X)=X^{p}$. According to [Kis06, §2.1.1], there exists an embedding $\mathfrak{S}\hookrightarrow\operatorname{W}(\mathcal{R})$ that preserves the Frobenius. Let $\mathscr{E}$ be the fraction field of the $p$-adic completion of $\mathfrak{S}[1/X]$. The above embedding extends to an embedding $\mathscr{E}\hookrightarrow\operatorname{W}(\operatorname{Frac}(\mathcal{R}))$, and we denote by $\mathscr{E}^{\mathrm{ur}}$ the maximal unramified extension of $\mathscr{E}$ inside $\operatorname{W}(\operatorname{Frac}(\mathcal{R}))[1/p]$, and by $\mathcal{O}_{\mathscr{E}^{\mathrm{ur}}}$ its ring of integers. We then define
$\mathfrak{S}^{\mathrm{ur}}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\mathcal{O}_{\mathscr{E}^{\mathrm{ur}}}\cap\operatorname{W}(\mathcal{R}).$
Since the absolute Galois group of $L_{\infty}$ fixes $\mathfrak{S}$, it acts naturally on $\mathfrak{S}^{\mathrm{ur}}$.
3.1.1. Breuil-Kisin modules
A Breuil-Kisin module $(\mathfrak{M},\varphi)$ is a finite free $\mathfrak{S}$-module $\mathfrak{M}$ equipped with a $\varphi_{\mathfrak{S}}$-semilinear Frobenius endomorphism $\varphi\colon\mathfrak{M}\rightarrow\mathfrak{M}$ whose cokernel of its $\mathfrak{S}$-linearization $1\otimes\varphi\colon\mathfrak{S}\otimes_{\mathfrak{S},\varphi}\mathfrak{M}\rightarrow\mathfrak{M}$ is killed by $E(X)$. Together with $\varphi$-equivariant $\mathfrak{S}$-linear maps, these objects form the category $\operatorname{\mathbf{BT}}_{\mathfrak{S}}(\varphi)$. Furthermore, Kisin obtained in [Kis06] a functor $\operatorname{\mathbf{D}}_{\mathfrak{S}}^{*}$ such that the following diagram commutes:
$\operatorname{\mathbf{Rep}}_{\mathbb{Q}_{p}}^{\mathrm{cris},\{0,1\}}(G_{L})$$\operatorname{\mathbf{MF}}^{\mathrm{adm},\{0,1\}}_{L}(\varphi)$$\operatorname{\mathbf{Rep}}_{\mathbb{Z}_{p}}^{\mathrm{cris},\{0,1\}}(G_{L})$$\operatorname{\mathbf{BT}}_{\mathfrak{S}}(\varphi)$$\scriptstyle{\operatorname{\mathbf{D}}^{*}_{\mathrm{cris}}}$$\scriptstyle{\operatorname{\mathbf{D}}_{\mathfrak{S}}^{*}}$$\scriptstyle{\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}-}$$\scriptstyle{D}$
where $D$ denotes the functor defined in [Kis06, §1], while $\{0,1\}$” stands for the weights, or the jumps, in $\{0,1\}$. The full subcategory of Breuil-Kisin modules corresponding to the representations of weight $1$ is denoted by $\operatorname{\mathbf{BT}}_{\mathfrak{S}}^{\mathrm{f}}(\varphi)$, and its objects are called formal Breuil-Kisin modules.
3.1.2. Construction
Set $M_{f}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{Hom}_{\mathbb{Z}_{p}[G_{L_{\infty}}]}(T_{f},\mathfrak{S}^{\mathrm{ur}})$ to be the free $\mathfrak{S}$-module of rank $1$ appearing in [Kis06, Lemma 2.1.10]. The action of the power series $f$ on $T_{f}$ extends to $M_{f}$ to be the multiplication by its linear coefficient. Indeed, given a generator $\pi_{f}\mapsto s$ of $M_{f}$, one has $f(\pi_{f}\mapsto s)=f^{\prime}(0)(\pi_{f}\mapsto s)$ because $f(\pi_{f})=f^{\prime}(0)\pi_{f}$. We do not loose any information on the remaining coefficients of $f$ because they are encoded in the action of $\mathbb{Z}_{p}\subset\mathfrak{S}$ on $M_{f}$. We let
$\mathfrak{M}_{f}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{\mathbf{D}}_{\mathfrak{S}}^{*}(T_{f})$
which is a rank-$1$ formal Breuil-Kisin module. By [Kis06, Lemma 2.1.10], we may assume that $\mathfrak{M}_{f}\subseteq M_{f}$ as $\mathfrak{S}$-modules. Consequently, the action of $f$ on $M_{f}$ restricts to $\mathfrak{M}_{f}$, and the inclusion guarantees that the latter encapsulates the other coefficients of the formal power series $f$.
3.2. The associated window
Let $S$ be the $p$-adic completion of the PD-envelope of $\operatorname{W}(\mathbb{F}_{q})[X]$ with respect to the ideal $(E(X))$, and let $\operatorname{Fil}^{1}S\subset S$ be the closure of the ideal generated by the divided powers of $E(X)$. As topological rings, we have an identification $S/\operatorname{Fil}^{1}S\simeq\mathcal{O}_{L}$. By construction, we have an inclusion with the Kisin ring $\mathfrak{S}\subset S$, and the Frobenius $\varphi$ on $\mathfrak{S}$ naturally extends to a $\sigma$-semilinear Frobenius endomorphism $\varphi$ on the ring $S$.
3.2.1. $S$-windows
Let $R\in\{\mathcal{O}_{L}/(p^{n})\,;\,n\geqslant 0\}$ be a ring. An $S$-window $(W,W^{\prime},\varphi)$ over $R$ is a finite projective $S$-module $W$ together with a distinguished submodule $W^{\prime}$ containing $(\operatorname{Fil}^{1}S)W$ such that $W/W^{\prime}$ is a projective module over $R$, and a $\sigma$-semilinear Frobenius endomorphism $\varphi\colon W\rightarrow W$ satisfying $\varphi(W^{\prime})\subseteq pW$; $W$ is generated as an $S$-module by $\varphi(W)+\varphi/p(W^{\prime})$. Given that $pW\subseteq(1\otimes\varphi)(\varphi^{*}(W))$, with $\varphi^{*}(W)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=S\otimes_{S,\varphi}W$, the map $1\otimes\varphi$ is injective and there exists a unique $S$-linear map $\psi\colon W\rightarrow\varphi^{*}(W)$ such that the composite $(1\otimes\varphi)\circ\psi\colon W\rightarrow W$ is the multiplication-by-$p$ map. We further require, for $n$ sufficiently large, that
$\varphi^{n-1*}(\psi)\circ\varphi^{n-2*}(\psi)\circ\cdots\circ\psi\colon W\longrightarrow\varphi^{n*}(W)$
to factor through $(p,\operatorname{Fil}^{1}S)\varphi^{n*}(W)$. Such objects form the category $S\textbf{-}\mathbf{Win}/R$, whose morphisms are $\varphi$-equivariant $S$-linear maps that preserve the distinguished submodule. In the particular case where $R=\mathcal{O}_{L}$, the data of an $S$-window over $\mathcal{O}_{L}$ is equivalent to
the data of a system $(W_{n})_{n\geqslant 1}$, where $W_{n}$ is an $S$-window over $\mathcal{O}_{L}/(p^{n})$ and is obtained from $W_{n+1}$ by base change for every integer $n\geqslant 1$. The extension
$S\otimes_{\varphi,\mathfrak{S}}-\colon\operatorname{\mathbf{BT}}_{\mathfrak{S}}^{\mathrm{f}}(\varphi)\longrightarrow S\textbf{-}\mathbf{Win}/\mathcal{O}_{L}$
defined in [Kis09, §1.2.3] is an equivalence of categories.
3.2.2. Construction
We set
$W_{f}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=S\otimes_{\varphi,\mathfrak{S}}\mathfrak{M}_{f}$
to be the $S$-window over $\mathcal{O}_{L}$ associated to our formal Breuil-Kisin module. Since Zink’s functors in [Zin01] for constructing $p$-divisible groups are formulated in terms of the limit description, we shall work with this perspective. Hence let $(W_{f,n})_{n\geqslant 1}$ be the corresponding system. Upon passing to the quotient of $W_{f}$ with which $W_{f,n}$ is identified, we may assume that each $W_{f,n}$ is a quotient of $W_{f}$. The action of $f$ on $\mathfrak{M}_{f}$ extends to an action of $f$ on $W_{f}$ by the multiplication-by-its-linear-coefficient map. This action descends compatibly to each quotient $W_{f,n}$. Taking the $W_{f,n}$ as quotients of $W_{f}$ guarantees that the remaining coefficients of $f$ are uniquely determined by the action of $\mathbb{Z}_{p}\subset S$ on each of these objects.
3.3. The associated $p$-divisible group
Let $(p\textbf{-}\mathbf{Div}/\mathcal{O}_{L})^{\mathrm{c}}$ be the category of finite-height connected $p$-divisible groups over $\mathcal{O}_{L}$. In [Zin01], Zink introduced two functors, $\operatorname{Dsp}$ and $\operatorname{BT}$, such that
$\operatorname{BT}\circ\operatorname{Dsp}\colon S\textbf{-}\mathbf{Win}/\mathcal{O}_{L}\longrightarrow(p\textbf{-}\mathbf{Div}/\mathcal{O}_{L})^{\mathrm{c}}$
is an equivalence. These functors are defined on $S$-windows over $\mathcal{O}_{L}$ expressed as a limit. Let $n\geqslant 1$ and write $R_{n}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\mathcal{O}_{L}/(p^{n})$. Set $P_{f,n}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{Dsp}(W_{f,n})=\operatorname{W}(R_{n})\otimes_{\varkappa_{n},S}W_{f,n}$ which is a $\operatorname{W}(R_{n})$-module, where $\varkappa_{n}\colon S\rightarrow\operatorname{W}(R_{n})$ is the map defined in [Zin01, p. 499]. The action of $f$ on $W_{f,n}$ extends to an action on $P_{f,n}$ and again is the multiplication-by-its-linear-coefficient map. All other coefficients of $f$ are uniquely determined by the action of $\mathbb{Z}_{p}\subset\operatorname{W}(R_{n})$ on $P_{f,n}$. Then, we set
[…]
which is a height-$1$ connected $p$-divisible group over $\mathcal{O}_{L}$. Denoting by $\operatorname{W}_{\mathrm{f}}$ the functor of Witt vectors with finite support, one defines $\Gamma_{f,n}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{BT}(P_{f,n})$ as a certain quotient of the functor $\operatorname{W}_{\mathrm{f}}(-)\otimes_{\operatorname{W}(R_{n})}P_{f,n}$. The action of $f$ on $P_{f,n}$ extends to the latter functor, descends to the quotient $\Gamma_{f,n}$ and is compatible with the transition maps. Consequently, the action of $f$ passes to the limit $\Gamma_{f}$ and is the multiplication-by-its-linear-coefficient map. The remaining coefficients of $f$ are uniquely determined by the action of $\mathbb{Z}_{p}$, that is, by the structural data of $\Gamma_{f}$ arising from the structure of $T_{f}$.
3.4. The main theorem
We have thus constructed $\Gamma_{f}$, a height-$1$ connected $p$-divisible group over $\mathcal{O}_{L}$, whose followed path is summarized below:
$\operatorname{\mathbf{Rep}}_{\mathbb{Z}_{p}}^{\mathrm{cris},1}(G_{L})$$\operatorname{\mathbf{BT}}^{\mathrm{f}}_{\mathfrak{S}}(\varphi)$$S\textbf{-}\mathbf{Win}/\mathcal{O}_{L}$$(p\textbf{-}\mathbf{Div}/\mathcal{O}_{L})^{\mathrm{c}}$$T_{f}$$\mathfrak{M}_{f}$$W_{f}$$\Gamma_{f}$}
and ensure that $f$ is an endomorphism of each of these objects. It follows that Lubin’s conjecture holds in our setting.
Theorem 3.1.
There exists a formal group $F$ over $\mathcal{O}_{K}\!$ such that $f,u\in\operatorname{End}_{\mathcal{O}_{K}}(F)$.
Proof.
The group law of $\Gamma_{f}$ induces a comultiplication on its Hopf algebra; this is a map $\mathcal{O}_{L}[\![T]\!]\rightarrow\mathcal{O}_{L}[\![X,Y]\!]$ given by sending $T$ to $F(X,Y)$, where $F(X,Y)$ is a height-$1$ formal group defined over $\mathcal{O}_{L}$. By construction, $f$ is an endomorphism of $F$, and so $F$ is defined over $\mathcal{O}_{L}\cap K=\mathcal{O}_{K}$ by the uniqueness property stated in §1.1. As $u$ is defined over $\mathcal{O}_{K}$ and commutes with $f$, it is necessarily an $\mathcal{O}_{K}$-endomorphism of $F$ as well. ∎